OpenStudy (anonymous):

1/9z=0 how do you get z alone and solve this.

jimthompson5910 (jim_thompson5910):

is the equation \[\large \frac{1}{9}z = 0\] OR is the equation \[\large \frac{1}{9z} = 0\]

OpenStudy (anonymous):

it is the top one

jimthompson5910 (jim_thompson5910):

Multiply both sides by 9 to get \[\large \frac{1}{9}z = 0\] \[\large 9*\frac{1}{9}z = 9*0\] \[\large 1z = 0\] \[\large z = 0\]

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np